OpenStudy (anonymous):

let g(x)= (x^2+x+2)/(x^(2/5)) calculate g'(x) without using the quotient rule

OpenStudy (lalaly):

g(x) = x^8/5 + x^3/5 + 2x^-2/5 g'(x)=8/5 x^3/5 + 3/5 x^-2/5 - 4/5 x^-3/5